A candy company sells bags of candy that are supposed to have an average of 40 pieces per bag, with = 9 . A random sample of 94 candy bags shows the average number of candies per box to be 43.1. Using a 1% level of significance, test that the average number of candies per bag is more than 40.

Determine the p-value and interpret the results.
Answer #1

µ = 40, x = 43.1, σ = 9, n = 94

Level of significance = 0.01

H0: µ = 40

H1: µ > 40

Test statistic: z = (x-µ)/(σ/n^0.5) = (43.1-40)/(9/94^0.5) = 3.34

p-value (Using Excel function NORM.S.DIST(z,cumultaive)) = NORM.S.DIST(3.34,TRUE) = 0.99958

Since p-value is more than 0.01, we do not reject the null hypothesis and conclude that µ = 40.

So, the average number of candies per bag is 40.